# An Expert is Worth One Token: Synergizing Multiple Expert LLMs as Generalist via Expert Token Routing

### v1

@@ 1 Introduction
In this work, we introduce a novel technique termed Expert-Token-Routing (ETR). ETR employs a hierarchical multi-expert LLM framework, wherein a meta LLM seamlessly orchestrates the allocation of tasks to expert LLMs. The core idea of ETR is encoding expert LLMs as special tokens within the vocabulary of the meta LLM. Under this formulation, the meta LLM can route to an expert LLM like generating new tokens. This approach allows for expert routing without training extra router model, simply by appending trained expert tokens into a frozen meta LLM. To train the expert tokens, we construct an automated pipeline to collect expert queries from existing datasets. RET can learn expert token embeddings from collected expert queries, which profiles the strengths of Expert LLMs, thereby handling situations where the expertise of expert LLMs is implicit or undisclosed.
@@
In our comprehensive experiments across six different domains, we compare ETR with the existing paradigms for expert LLM collaboration. Our findings indicate that PET significantly outperforms the baselines as a generalist, with an overall performance improvement of 5.64%. It is worth noting that our framework also integrates the advantages of existing paradigms, as illustrated in Table 1.
@@ 2.2 Learning Expert Embeddings via Expert Query Set
This embedding matrix $W_{E}$ are the only tunable parameters. Optimizing this parameter doesn’t require gradient propagation through the main part of the LLM parameters (Hao et al., 2023), leading to training that is both more stable and efficient than prompt tuning (Lester et al., 2021) or prefix tuning (Li and Liang, 2021). Thus, the GPU memory usage for expert token embedding tuning is nearly the same to that of LLM inference.

### v2

@@ 1 Introduction
In this work, we introduce a novel technique termed Expert-Token-Routing (ETR). ETR employs a multi-expert LLM collaboration framework, wherein a meta LLM can seamlessly switch to a specific expert LLM. The core idea of ETR is encoding expert LLMs as special tokens within the vocabulary of the meta LLM. Inspired by ToolkenGPT (Hao et al., 2023), which captures tool semantics using special tokens, ETR employs expert tokens to characterize the expertise of different expert LLMs. Under this formulation, the meta LLM can route to an expert LLM like generating new tokens. This approach allows for expert routing without training extra router model, simply by appending trained expert tokens into a frozen meta LLM. To train the expert tokens, we construct an automated pipeline to collect expert queries from existing datasets. ETR can learn expert token embeddings from collected expert queries, which profiles the strengths of expert LLMs, thereby handling situations where the expertise of expert LLMs is implicit or undisclosed.
@@
In our comprehensive experiments across six different domains, we compare ETR with the existing paradigms for expert LLM collaboration. Our findings indicate that ETR significantly outperforms the baselines as a generalist, with an overall performance improvement of 5.64%. It is worth noting that our framework also integrates the advantages of existing paradigms, as illustrated in Table 1.
@@ 2.2 Learning Expert Embeddings via Expert Query Set
This embedding matrix $W_{E}$ are the only tunable parameters. Therefore, similar to pioneering work on expanding the language modeling head of LLMs (such as ToolkenGPT (Hao et al., 2023)), training ETR doesn’t require gradient propagation through the main part of the LLM parameters, leading to training that is both more stable and efficient than prompt tuning (Lester et al., 2021) or prefix tuning (Li and Liang, 2021). Thus, the GPU memory usage for expert token embedding tuning is nearly the same to that of LLM inference.